OpenStudy (anonymous):

multipy (4+ square root of 10)(4- square root of 10

OpenStudy (anonymous):

You have to foil 4 * 4 = 16 4 * -sqrt(10) = -4sqrt(10 sqrt(10) * 4 = 4 sqrt(10) sqrt(10) * - sqrt(10) = -10 Add this up.. 16 - 10 = 6 -4sqrt(10) + 4sqrt(10) = 0 So the answer = 6

OpenStudy (anonymous):

do you wanna help with mine?